• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Abdurahim

    Log entry time 13:09:20 on June15,2010

    Entry number 326737

    keyword=Compton Cavity has been restored

    Due to EPICS control for QWP was stuck, cavity is not locking. I did an EPICS reboot and power cycled the Suruga controller and then very thing came back. Just for future reference here are the parameters,

    === 06-15-2010-13-01 ===

    M1 (h, v) : 9695[c] 1395[c]
    M2 (h, v) : 104510[c] -34245[c]
    M3 (v) : 280000[c]
    CavityStatus : Cavity is locked
    IPGPower (W): 4.998
    IPG Temp in (C): 43.1
    IPG Temp out (C): 33.96
    IPG Diode Current:2.245
    GreenPower (mW): 1547.75
    PPLN Temp (C) : 60.76
    Lightwave Temp(C):+47.125
    Lightwave Power: Pwr 79mW
    RampAmplitude(C): 0.08
    RampPeriod (s): 180
    Central Temp (C): 47.201
    Q1(x, y, s) : -124.711 635.036 -0.435818
    Q2(x, y, s) : 822.571 -21.0068 -0.841116
    PDR (V) : 0.18861
    PDT (V) : 0.778246
    Servo Gain : 125
    TrackingLevel (V):0.55
    Cavity Power (W): 2688.15
    S1 power (mW): 5.86642
    S2 power (mW): 398.643
    Polarization (%): 97.0854
    PolarizationState:LEFT

    ===



    FIGURE 1